OpenStudy (anonymous):

find the solution set of each of the following linear equations. a) 7x -5y = 3 b) -8x1 + 2x2 - 5x3 + 6x4 = 1

OpenStudy (turingtest):

a) solving for x we get\[x=\frac{3}{7}+\frac{5}{7}t\]\[y=\left\{ t|t \in R \right\}\]

OpenStudy (turingtest):

b) solving the next for x1 we get\[x_1=\frac{1}{4}r-\frac{5}{8}s+\frac{3}{4}t-\frac{1}{8}\]\[x_2=r\]\[x_3=s\]\[x_4=t\]\[r,s,t\in R\]now what's so hard about that???

OpenStudy (turingtest):

a) may be better stated this way..\[x=\frac{3}{7}+\frac{5}{7}t\]\[y=t\]\[t\in R\]same idea though